Given a firms demand function Q-90+2P=0 and it's average cost function AC=Q²-8Q+57+2/Q. Find the level of output which maximises marginal cost.


1
Expert's answer
2021-05-25T16:55:43-0400

Given Information

Demand function

"Q-90+2P=0"

Average cost function

"AC=Q²-8Q+57+\frac{2}{Q}" 

"Total\space cost = AC \times quantity"

        "=( Q²-8Q+57+\frac{2}{Q} ) \times Q"

"=Q^{3} - 8Q^{2} +57Q + 2"

      Differentiate Total Cost with respect of Quantity

"MC = \frac{\delta TC}{\delta Q}"

"MC = \frac{Q^{3} − 8Q^{2} +57Q + 2}{Q}"

"= 3Q^{2} -16Q +57"

To calculate Maximization of MC of output we equate Mc with 0

MC = 0

"3Q^{2} -16Q +57 = 0"

a = 3

b = -16

c = 57

"b^{2} - 4ac = (-16)2 - 4 \times 3\times 57"

"= 256 - 684"

"= -428."

"\frac{−b ±√b2 − 4ac}{2a}"


 "= \frac{−(−16) ±√ −428}{2 x 3}"

"= 36.68i"

And another value will be -4.68i

 Negative output is not possible hence the 36.68 output will be maximize MC.